A 0.060-kg tennis ball, moving with a speed of 5.46 m/s , has a head-on collision with a 0.090-kg ball initially moving in the same direction at a speed of 3.54 m/s . Assume that the collision is perfectly elastic.

a) A 0.060-kg tennis ball, moving with a speed of 5.46 m/s , has a head-on collision with a 0.090-kg ball initially moving in the same direction at a speed of 3.54 m/s . Assume that the collision is perfectly elastic.

b) Determine the direction of the velocity of the 0.060- ball after the collision.

in the direction of the initial velocity
in the direction opposite to the initial velocity

c) Determine the speed of the 0.090-kg ball after the collision.

d)  Determine the direction of the velocity of the 0.090- ball after the collision.

in the direction of the initial velocity
in the direction opposite to the initial velocity
Answer #1

answer) we have to apply the conservation of momentum, so we have

m1u1+m2u2=mu+mv

0.06*5.46+0.09*3.54=0.06*u+0.09*v

0.6462=0.06u+0.09v..................1)

now from conservation of energy we have

v-u=u1-u2

v-u=5.46-3.54=1.92

v=1.92+u................2)

replacing the value of v in eqn 1)

0.6462=0.06u+0.09(1.92+u)

0.6462=0.06u+.09u+0.1728

0.4734=0.15u

u=3.16 m/s

answer to part a) 3.16 m/s or 3.20 m/s

part b) since the speed is positive the answer is =in the direction of initial velocity

part c) using eqn 2) we have

v=1.92+3.16=5.08 m/s

so the answer is 5.08 m/s or 5.10 m/s

part d) since the speed is positive the answer= in the direction of initial velocity